OpenStudy (anonymous):

Evaluate the series 50 + 10 + 2 + . . .

OpenStudy (anonymous):

I'm confused as to what "Evaluate" actually means. Answer choices: A) 62 B) 1/5 C) The series diverges; it does not have a sum D) 62.5

OpenStudy (solomonzelman):

you divide by 5 every time. you start from 50 so you can write it as \(\large\color{black}{ \displaystyle \sum_{ n=0 }^{ \infty } ~ \dfrac{50}{5^n}}\)

OpenStudy (solomonzelman):

this series definitely converges, will you give me that?

OpenStudy (anonymous):

Yeah

OpenStudy (solomonzelman):

you can find the exact sum because it is geometric...

OpenStudy (solomonzelman):

\(\Large\displaystyle\sum_{ n=0 }^{ \infty } ~ 50(1/5)^n=\frac{50}{1-\frac{1}{5}}=?\)

OpenStudy (anonymous):

we would get 125/2 or 62.5. Thanks. Is there any place where i could get a quick refresher on Geometric and arithmetic series?

OpenStudy (anonymous):

That is correct